What is the oxidation number for each element in the given compound

a) SO2
b) Xe
c)CO
d)K2Cr2O7
e)NO3-
1
Expert's answer
2020-11-12T06:37:57-0500

a) SO2 . Oxidation for Oxygen = -2

S+2(-2) = 0

S -4 = 0

S =4 . Oxidation number of sulphur is 4

b) Xe Oxidation number of xenon = 0

c) CO Oxidation number of oxygen = -2

C-2=0. C= 2 . Oxidation number of carbon is 2

d) K2Cr2O7. Oxidation number of potassium = 1

Oxidation number of oxygen = -2

2(1) +2Cr +7(-2) = 0

2 + 2Cr -14 =0

2Cr -12= 0

2Cr = 12

Cr = 12/2 = 6 . Oxidation number of chromium= 6

e) NO3- Oxidation number of oxygen is -2

N+3(-2) = -1

N-6 =-1

N= -1+6 =5 Oxidation number of Nitrogen =5